OpenStudy (anonymous):

Factor completely −3x2 + 6x – 9.

myininaya (myininaya):

well first do you see a common factor amongst -3,6,-9?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@AmberLOVESyou!

OpenStudy (anonymous):

6

myininaya (myininaya):

so -3=-3 6=2(3) -9=-3(3) you don't see a common factor amongst -3,6, and -9?

OpenStudy (anonymous):

3

myininaya (myininaya):

yes or you can also say -3 either way

OpenStudy (anonymous):

yea but those are not my choices

myininaya (myininaya):

\[-3(x^2-2x+3)\]

myininaya (myininaya):

now try to see if you can factor x^2-2x+3

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

that's the answer ?

myininaya (myininaya):

what two numbers multiply to be 3 and add up to be -2

myininaya (myininaya):

that is right because there are no two factors of 3 that have product 3 and sum -2

OpenStudy (anonymous):

thank you :)